OpenStudy (salina):

Find counterexample for statement 2^2 +3^2 is divisible by 4 the key n=3

myininaya (myininaya):

4+9=13 13 isn't divisible by 4

OpenStudy (anonymous):

Well 4 can't divide 4 +9 =13

OpenStudy (salina):

sorry \[2^n+3^n\] not 2^2+3^2

OpenStudy (anonymous):

Ohhh I see

myininaya (myininaya):

i cant read both have same exponent right?

OpenStudy (anonymous):

ahun to the power of n

OpenStudy (salina):

yes,

myininaya (myininaya):

2^3+3^3=8+9=17 4 doesnt divide 17 this is one counterexample

OpenStudy (salina):

the key is n=3 divi by 4

OpenStudy (anonymous):

myininaya what is counterexample ?

myininaya (myininaya):

giving an example of showing a statement is false

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

2^3 + 3^3 = 35

myininaya (myininaya):

2^1+3^1=2+3=5 5 is not divisible by 4 this is another counterexample

OpenStudy (anonymous):

yeah which is still not divisible by 4

myininaya (myininaya):

what do you mean the key n=3?

OpenStudy (salina):

the question has key answer n= 3

OpenStudy (anonymous):

You Put anything instead of n but you won't get it divisible by 4 ...maybe that's what I think ...the problem lies in three \[\frac{(2 \times 2 \times 2 \times ...n) + ( 3 \times 3 \times .....n)}{4} \]

myininaya (myininaya):

there are other n's like n=1

myininaya (myininaya):

and n=2

myininaya (myininaya):

2^4+3^4 =16+81 =97 97 isn't divisible 4 do you have enough counterexamples? lol

OpenStudy (salina):

I no idea take me for ever

myininaya (myininaya):

do you have any questions?

OpenStudy (salina):

no.I don't